# Comprehensive Review of Deep Unfolding Techniques for Next-Generation Wireless Communication Systems

### v2

@@ III Deep Unfolding: General Structure
Figure 4: General framework of deep unfolding.
For any iterative algorithm, the update equation for each iteration is given as:
${\bf{s}}^{(t+1)}=h({\bf{s}}^{t};{\bf{\Psi}}^{t})$
where ${\bf{s}}^{t}\in{\mathbb{R}}^{n}$ denotes the variable vector to be updated in each iteration in the range $t=1,\dots,T$. The iterative function for the iterative algorithm is represented by $h(\cdot;\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. The trainable parameters, ${\bf{\Psi}}^{t}\in\mathbb{R}^{n}$ include the model parameters and the step size for the regularization coefficients of the algorithm [25]. Each iteration function $h(\cdot;\cdot)$ is mapped to a single network layer, and a finite number of layers are stacked together as shown in Fig. 4. To learn the trainable parameters ${\bf{\Psi}}^{t}$ in an end-to-end approach, the following optimization is carried out:
$\text{minimize}_{{\bf{\Psi}}_{T}}\mathcal{L}({\bf{s}}^{T}({\bf{\Psi}}_{T}))$
@@
where $\mathcal{L}(\cdot)$ denotes the loss function for training, ${\bf{\Psi}}_{T}=\{\Psi^{t}\}_{t=0}^{T}$ is the set of trainable parameters of the total $T$-layer network and ${\bf{s}}^{T}(\cdot)$ is the output of the unfolded network. Feeding the data forward through the unfolded network with trainable parameters in the testing stage is similar to executing the parameter-optimized iterative algorithm for a finite number of iterations. Thus, the number of layers in deep unfolding models is fixed, resulting in a predefined computational complexity. Within these fixed layers, the model is trained to optimize the trainable parameters, enabling the problem to be addressed using a structured algorithm with parameters tailored for optimal performance.
The procedure to solve an optimization problem using a deep unfolding approach is outlined below.
  1. 1.
    Identify a principle algorithm that serves as the foundation for deep unfolding.
  2. 2.
    Select trainable parameters or specific blocks in the principle algorithm. Replace these with data-trained operations optimized through deep unfolding.
  3. 3.
    Define the loss function and represent the algorithm as a DNN. Each layer corresponds to an iteration of the original algorithm. Train the network using supervised or unsupervised methods. Since the number of iterations is predetermined, the network’s layers match these iterations. The number of layers can also be fine-tuned to balance performance and computational complexity.
TABLE III: Key technologies envisioned for 6G using deep unfolding.
Key technology Task Reference
Massive MIMO Detection [49],[50], [51], [52], [53], [44], [54], [55], [56], [57], [58], [59], [60], [61]
Channel estimation [62], [63], [64], [65], [66], [67]
Precoder Design [68], [69], [70], [71], [72], [73], [74], [75], [76], [77], [78], [79], [80], [81]
ISAC [82], [83], [84]
mmWave Channel Estimation [85]
Precoder design [86], [87], [88]
ISAC [89]
Modern Multicarrier Modulation OFDM Detection [90],[91], [92]
OTFS Detection [46]
OFDM Channel Estimation [93], [94], [95], [45], [96], [97], [98], [99], [47] [100]
OTFS Channel Estimation [101]
OFDM-ISAC [102], [103]
OTFS-ISAC [104]
Next Generation Multiple Access Detection [105], [106]
Precoder design [107]
Orbital Angular Momentum Detection [108]
IRS Channel Estimation [109], [110], [111], [112]
Precoder design [113], [114], [115], [116]
ISAC [117]
THz Precoder design [118], [119], [120]
Principle algorithms are typically iterative and based on optimization techniques. However, they are often complex, computationally demanding, and slow to converge. Deep unfolding improves these algorithms by leveraging data training. The key contributions of deep unfolding are:
  1. 1.
    Optimizing hyperparameters to enhance convergence, leading to better objective values or faster processing [26, 76, 84, 11].
  2. 2.
    Simplifying complex tasks, such as matrix inversion or decomposition, with lower computational complexity [121, 70, 122].
The first application is common in unfolded projected gradient ascent/descent (PGA/PGD) methods, while the second is widely used in unfolded weighted minimum mean square error (WMMSE) methods. Another approach involves learning the output at each layer, which corresponds to the iterative solution of the principal algorithm. This technique is widely applied in detection networks [123, 118, 49], following the PGA/PGD framework.
The main advantage of deep unfolding is its ability to incorporate domain knowledge into the network architecture. This improves performance while reducing complexity by using fewer parameters and a fixed number of layers. As a result, deep unfolding requires a smaller training dataset compared to generic DNNs. Next, we provide a comprehensive survey of deep unfolding algorithms designed for various signal processing tasks in wireless communication. Additionally, TABLE III lists key 6G technologies relevant to deep unfolding-based signal processing.
IV Signal Detection

### v3

@@ II Deep Unfolding: General Architecture and Fundamentals
Figure 2: General framework of training for deep unfolding. The learnable parameters are shown in red color.
@@
To learn the trainable parameters $\left\{\boldsymbol{\theta}\right\}_{l=1}^{L}$ in an end-to-end approach, the loss $\ell({\bf{s}}^{L},\bf{s})$ is computed, where ${\bf{s}}^{L}$ is the output of the network and $\bf{s}$ is the ground truth target. The key idea behind end-to-end learning is that all parameters across $L$ layers are simultaneously adjusted through backpropagation, where the loss is minimized using gradient-based optimization. Feeding the data forward through the unfolded network with trainable parameters in the testing stage is similar to executing the parameter-optimized iterative algorithm for a finite number of iterations. Thus, the number of layers in deep unfolding models is fixed, resulting in a predefined computational complexity. Within these fixed layers, the model is trained to optimize the trainable parameters, enabling the problem to be addressed using a structured algorithm with parameters tailored for optimal performance.
Example 1.
To develop an understanding on deep unfolding, we take ADMM as an example. ADMM decomposes a complex optimization problem into simpler subproblems, which are solved iteratively. The algorithm combines the principles of dual ascent and the method of multipliers. Each iteration involves solving the subproblems for the primal variables in an alternating manner, followed by updating the dual variables (Lagrange multipliers) to enforce consistency with the original problem’s constraints [37]. This iterative process ensures that the subproblem solutions converge to the global solution of the main optimization problem. ADMM in its unfolded version has been widely used in MIMO signal detection [25], [38] and decoding in error correction codes [39, 40]. Here, we show the unfolding of ADMM for MIMO detection as an illustrative example.
Figure 3: Deep unfolding training framework for ADMM-based MIMO detection. The learnable parameters are shown in red color.
Consider a MIMO system with $M$ transmit antennas and $N$ receive antennas. The received signal vector at the receiver can be given by
${\bf{\bar{y}}}={\bf{\bar{H}\bar{x}}}+{\bf{\bar{n}}}$
where ${\bf{\bar{H}}}\in\mathbb{C}^{N\times M}$ represents the channel matrix, ${\bf{\bar{x}}}\in{\cal{\bar{X}}}^{M}$ is the transmitted symbol vector, and ${\bf{\bar{n}}}\in\mathbb{C}^{N}$ denotes additive white Gaussian noise (AWGN) vector at the receiver. Here, $\cal{\bar{X}}$ represents the complex constellation set. In this example, we consider the QPSK modulation scheme, and thus, $\left({{\cal{\bar{X}}}=\left\{{1+i,-1+i,-1-i,1-i}\right\}}\right)$. The maximum likelihood (ML) detection rule is given as
$\min_{\bf{\bar{x}}\in{\cal{\bar{X}}}^{M}}\left\lVert{\bf{\bar{y}}}-{\bf{\bar{H}\bar{x}}}\right\rVert_{2}^{2}$ (2)
Denote the following real-valued vectors and matrix:
$\mathbf{y}\triangleq\begin{bmatrix}\mathfrak{R}(\bar{\mathbf{y}})\\ \mathfrak{I}(\bar{\mathbf{y}})\end{bmatrix},\quad\mathbf{x}\triangleq\begin{bmatrix}\mathfrak{R}(\bar{\mathbf{x}})\\ \mathfrak{I}(\bar{\mathbf{x}})\end{bmatrix},\quad\mathbf{n}\triangleq\begin{bmatrix}\mathfrak{R}(\bar{\mathbf{n}})\\ \mathfrak{I}(\bar{\mathbf{n}})\end{bmatrix},$
[…]
where $\mathfrak{R}(\cdot)$ and $\mathfrak{I}(\cdot)$ denote the real and imaginary part of a complex vector or matrix, respectively. With this transformation, each element of $\mathbf{x}$ is drawn from the discrete space {-1, 1} . We then relax this space to the continuous domain ${\cal{X}}=[-1,1]$. Consequently, (2) can be relaxed to the following problem in the real domain: $\min_{{\bf{x}}\in{{\cal{X}}}^{2M}}\left\lVert{\bf{y}}-{\bf{Hx}}\right\rVert_{2}^{2}.$ We further introduce an auxiliary variable $\mathbf{z}$ to reformulate this problem as:
$\underset{\begin{subarray}{c}{\bf{x}}\in{{\cal{X}}}^{2M}\end{subarray}}{\min}$ $\quad\left\lVert{\bf{y}}-{\bf{Hx}}\right\rVert_{2}^{2}$
subject to $\quad\mathbf{x}=\mathbf{z},\mathbf{z}\in{{\cal{X}}}^{2M}.$
The augmented Lagrangian is given as:
$\mathcal{L}_{\rho}(\mathbf{x},\mathbf{z},\boldsymbol{\lambda})=\left\lVert\mathbf{y}-\mathbf{Hx}\right\rVert_{2}^{2}+\boldsymbol{\lambda}^{T}(\mathbf{x}-\mathbf{z})\\ +\frac{\rho}{2}\left\|\mathbf{x}-\mathbf{z}\right\|_{2}^{2}$ (4)
where $\boldsymbol{\lambda}\in\mathbb{R}^{2M}$ is the dual variable (or Lagrange multiplier) and $\rho>0$ is the penalty parameter. The iterative steps of the ADMM involve alternating minimization of the augmented Lagrangian with respect to the primal variables $\bf x$ and $\bf z$, followed by an update of the dual variable, $\boldsymbol{\lambda}$ [41, 42]. The closed-form solutions for $\bf x$ and $\bf z$ can be obtained as:
${\mathbf{x}}^{l+1}$ $=\arg\min_{\mathbf{x}}\mathcal{L}_{\rho}(\mathbf{x},\mathbf{z}^{l},\boldsymbol{\lambda}^{l})$
$=\arg\min_{\bf{x}}\left\lVert\mathbf{y}-\mathbf{Hx}\right\rVert_{2}^{2}+{(\boldsymbol{\lambda}^{l})}^{T}({\mathbf{x}}-{\mathbf{z}}^{l})+\frac{\rho}{2}\left\lVert{\mathbf{x}}-{\mathbf{z}}^{l}\right\rVert_{2}^{2}$
$=\left({\mathbf{H}}^{T}\mathbf{H}+\rho\mathbf{I}\right)^{-1}\left({\mathbf{H}}^{T}\mathbf{y}+\rho{\mathbf{z}}^{l}-{\boldsymbol{\lambda}}^{l}\right),$ (5)
${\mathbf{z}}^{l+1}$ $=\arg\min_{\mathbf{z}}\mathcal{L}_{\rho}(\mathbf{x}^{l+1},\mathbf{z},{\boldsymbol{\lambda}}^{l})$
$=\arg\min_{\bf{z}}-{(\boldsymbol{\lambda}}^{l})^{T}{\mathbf{z}}+\frac{\rho}{2}\left\lVert{\mathbf{x}}^{l+1}-{\mathbf{z}}\right\rVert_{2}^{2}$
$=\arg\min_{\bf{z}}\frac{\rho}{2}\left\lVert{\mathbf{z}}-({\mathbf{x}}^{l+1}+\frac{{\boldsymbol{\lambda}}^{l}}{\rho})\right\rVert_{2}^{2}$
$=\Pi_{[-1,1]}\left(\mathbf{x}^{l+1}+\frac{\boldsymbol{\lambda}^{l}}{\rho}\right),\quad\quad{\text{and}}$ (6)
${\boldsymbol{\lambda}}^{l+1}$ $={\boldsymbol{\lambda}}^{l}+\rho\left({\mathbf{x}}^{l+1}-{\mathbf{z}}^{l+1}\right)\quad\forall l$ (7)
where $\Pi_{[-1,1]}({\boldsymbol{a}})$ denotes an element-wise projection of $\boldsymbol{a}$ onto the set $[-1,1]$. More specifically, if ${\boldsymbol{b}}=\Pi_{[-1,1]}(\boldsymbol{a})$, we have
$b_{i}=\begin{cases}+1,&\text{if }a_{i}>1\\ -1,&\text{if }a_{i}<1\\ a_{i},&\text{otherwise}.\end{cases}$
The detailed procedure for the unfolded ADMM approach is provided in Algorithm 1. In deep unfolding, to ensure gradient flow during backpropagation in the training stage, all operations must be differentiable. To enable differentiable approximation, a soft projection is used that allows gradient to flow while keeping outputs near the desired constellation points. Consequently, (6) becomes:
${\mathbf{z}}^{l+1}$ $=\Pi_{\text{soft}}\left(\mathbf{x}^{l+1}+\frac{\boldsymbol{\lambda}^{l}}{\rho}\right).$ (8)
Furthermore, the layer-wise unfolding of ADMM for MIMO detection is illustrated in Fig. 3. The received symbol vector $\mathbf{y}$ and the channel matrix $\mathbf{H}$ are treated as fixed inputs and are provided to every layer of the network. Each layer of the unfolded network corresponds to one iteration of the ADMM procedure, specifically the update steps in equations (5), (7), and (8). Corresponding to each of these equations, a specific $\rho$ is considered to enable learning. These parameters form the trainable set ${\boldsymbol{\theta}}=\left\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{1}^{(l)}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{2}^{(l)}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{3}^{(l)}}\right\}_{l=1}^{L}$, where each layer $l$ has its own set of penalty coefficients. In the final layer $L$, the network produces the estimate ${\mathbf{z}}^{L}=\hat{\bf{z}}$. For $n=1,\dots,N$, data samples, the output $\hat{\bf{z}}_{n}({\bf{y}}_{n},{\bf{H}};\rho_{1}^{(l)},\rho_{2}^{(l)},\rho_{3}^{(l)})$ is retrieved from the network as the predicted transmitted symbol for ${\bf{y}}$. The network training loss function formed between the predicted output $\hat{\bf z}$, and ground truth transmitted symbol ${\bf z}_{n}$ is given as:
$\ell(\rho_{1}^{(l)},\rho_{2}^{(l)},\rho_{3}^{(l)})=\frac{1}{N}\sum_{n=1}^{N}\left\lVert\hat{\bf{z}}_{n}({\bf{y}}_{n},{\bf{H}};\rho_{1}^{(l)},\rho_{2}^{(l)},\rho_{3}^{(l)})-{\bf z}_{n}\right\rVert_{2}^{2}.$ (9)
The network is trained by minimizing the loss with gradient-based learning techniques such as stochastic gradient descent (SGD) to learn $\rho_{1}^{(l)}$, $\rho_{2}^{(l)}$ and $\rho_{3}^{(l)}$.
Algorithm 1 Unfolded ADMM for MIMO detection.
1:  Input: Initial values $\mathbf{z}^{(0)}$, ${\boldsymbol{\lambda}}^{(0)}$
2:  Given: $\mathbf{y}$ and $\mathbf{H}$
3:  Learnable parameters $\boldsymbol{\theta}=\left\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{1}^{(l)}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{2}^{(l)}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{3}^{(l)}}\right\}_{l=1}^{L}$
4:  for $l=1$ to $L$ do
5:     $\mathbf{x}^{l+1}=\left(\mathbf{H}^{T}\mathbf{H}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{1}^{(l)}}\mathbf{I}\right)^{-1}\left(\mathbf{H}^{T}\mathbf{y}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{1}^{(l)}}\mathbf{z}^{l-1}-{\boldsymbol{\lambda}}^{l}\right)$
6:     $\mathbf{z}^{l+1}=\Pi_{\text{soft}}\left(\mathbf{x}^{l+1}+\frac{{\boldsymbol{\lambda}}^{l-1}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\rho_{2}^{(l)}}}}\right)$
7:     ${\boldsymbol{\lambda}}^{l+1}={\boldsymbol{\lambda}}^{l}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{3}^{(l)}}\left({\mathbf{x}}^{l+1}-{\mathbf{z}}^{l+1}\right)$
8:  end for
9:  Output: Final estimate $\hat{\bf{z}}={\bf{z}}^{L}$
[…]
Figure 4: BER performance comparison for 16$\times$16 MIMO system. For soft projection, $tanh$ is used.
Fig. 4 presents the bit error rate (BER) performance comparison of the unfolded ADMM detector with zero forcing (ZF), minimum mean squared error (MMSE), ADMM and sphere decoder (SD), under QPSK modulation. We consider a $16\times 16$ MIMO system with Rayleigh fading channels. The results demonstrate that the unfolded ADMM detector outperforms ZF, MMSE, and standard ADMM detectors, except for SD. The unfolded ADMM detector performs better due to the use of deep unfolding, where learnable parameters are adaptively trained. This example has been included to provide the readers with an illustrative understanding of the deep unfolding process. The comparative analysis highlights the advantage of the unfolded method over classical approaches. Additional performance improvements can be achieved by optimizing or introducing trainable parameters in the unfolded network.
In the following sections, we provide a comprehensive survey of deep unfolding algorithms for a variety of critical wireless communication tasks.
III Signal Detection